# FCOC: A Fractal-Chaotic Co-driven Framework for Financial Volatility Forecasting

### v1

@@ 1 Introduction
The forecasting of financial market volatility, as the second moment of asset returns, is a cornerstone of modern finance, crucial for risk management, option pricing, and portfolio allocation [1]. While deep learning methods have shown immense potential in this domain due to their powerful nonlinear pattern recognition capabilities [2, 3], their full potential remains constrained by two fundamental and often-overlooked bottlenecks in the standard modeling pipeline.
The first bottleneck is feature fidelity. Financial time series, as typical complex systems, possess an intrinsic structure that traditional statistical features cannot fully capture. The Fractal Market Hypothesis (FMH) from econophysics [4] points out that long-range memory and multifractal properties are key dynamical characteristics of the market. This makes fractal analysis a powerful and theoretically grounded tool for our investigation. However, the application of these advanced tools faces significant limitations. First, their analytical framework is often limited to a single time series, largely failing to consider the cross-asset risk transmission mechanisms, thus ignoring the asymmetric cross-correlation phenomenon that is at the core of systemic risk [5, 6, 7]. Second, the standard implementation of these methods has inherent instability, is prone to introducing spurious noise, and thus harms the fidelity of the extracted features [8]. These two problems together lead to a situation where the input information the model relies on is, at its source, a distorted representation of the market’s true state.
The second, and more fundamental, bottleneck lies in model responsiveness. The core processing units widely adopted in deep learning models—static activation functions such as ReLU—reveal their inherent design limitations when processing highly dynamic and non-stationary financial signals. This limitation stems from a complexity mismatch [9]: a simple, fixed-logic processor is used to analyze an essentially chaotic and dynamic complex signal. This bottleneck means that even with perfect input features, the model’s rigid internal processing mechanism cannot provide an effective dynamic response, ultimately leading to information decay and loss during internal transmission [10].
@@ 2.1 Deep Learning Models for Financial Time Series Forecasting
Modern financial forecasting methods have widely adopted deep learning architectures. Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) [2, 3] and Gated Recurrent Units (GRU) [11], are commonly used for their ability to handle temporal dependencies. In recent years, more advanced architectures such as the Transformer [12] and state-space models like Mamba have also been explored. For instance, Lu and Xu [13] proposed an efficient Time-series Recurrent Neural Network (TRNN) to improve training efficiency. Chen et al. [14] developed a hybrid model combining a Temporal Convolutional Network (TCN) with BiGRU for new energy stock index forecasting. While these works have made significant progress in model architecture, they almost universally rely on traditional static activation functions, thus failing to fundamentally address the internal model responsiveness problem.
@@ 2.2 Multifractal Analysis of Financial Time Series
To overcome the informational limitations of raw time series, a more theoretically profound approach originates from econophysics. This discipline aims to uncover the intrinsic physical laws of the market, with the multifractal nature of financial time series being a key insight that reveals the non-uniformity of volatility across different time scales [15].
In response, econophysics has developed an evolving toolkit of methods, progressing from Detrended Fluctuation Analysis (DFA) [4] to Multifractal DFA (MF-DFA) [15]. A significant advance came with the recognition of market asymmetry, leading to Asymmetric MF-DFA (A-MFDFA) [16, 17]. However, these methods were long confined to single-asset analysis, overlooking the systemic interactions of the market as a whole. In fact, asymmetric cross-correlations between different assets are a key driver of systemic market risk, especially during downturns [5, 6, 7]. Recent work has continued to validate this, with Yu et al. [18] successfully combining multifractal analysis with GRU networks and Wang and Lee [19] applying modified MF-ADCCA to forecast stock index volatility. Although subsequent research began to address cross-correlations with more complex techniques, they commonly inherited a persistent technical flaw from earlier methods: the use of non-overlapping segmentation, which has been shown to introduce spurious fluctuations and compromise the stability of fractal measurements [8]. Our Fractal Feature Corrector (FFC) is designed to solve both of these historical problems simultaneously. By using Multifractal Asymmetric Detrended Cross-Correlation Analysis (MF-ADCCA) [20] as its core to capture systemic interactions, and employing a robust Overlapping Sliding Window (OSW) implementation to correct for instability [21], the FFC aims to provide a more reliable and higher-fidelity set of complexity features than previously available.
@@ 2.3 Dynamic and Chaotic Activation Functions
The concept of enhancing neural networks with internal dynamic components originates from computational neuroscience. Unlike standard neural networks, the biological brain is believed to operate based on continuous and chaotic neural oscillations [22]. Inspired by this, academia has begun to explore replacing static activation functions with dynamic systems [9]. This direction also aligns with cutting-edge paradigms in econophysics, which posit that financial markets can be viewed as a dynamic energy field where assets behave like interacting, chaotic oscillators. From this perspective, an oscillator with excitatory and inhibitory mechanisms is not merely a computational tool but a physical analogy for the fundamental push-pull dynamics of market forces.
Therefore, our adoption of the Chaotic Oscillation Component (COC) is dually motivated: computationally, it solves the complexity mismatch problem; physically, it provides a more plausible model for the phenomena we aim to forecast. Our COC is built upon the foundational work on the Lee oscillator [23]. While generative models like the VAR-VAE proposed by Leushuis [24] have started to incorporate probabilistic dynamics in latent spaces, our work focuses on a different, complementary goal: embedding dynamics directly into the neuron’s fundamental activation process. Our core contribution in this area is not only the systematic application of this concept to financial volatility forecasting but also the engineering of two novel oscillator configurations (T9 and T10) through parameter space exploration, specifically to better capture the unique dynamics of financial markets, such as abrupt regime switching.
@@ 3.1 Fractal Feature Corrector (FFC)
This process yields $N_{seg}=\lfloor(N-s)/s_{step}\rfloor+1$ overlapping segments. In this work, an overlap ratio of $\rho=1/3$ is used, a choice that strikes a balance between enhancing statistical stability and maintaining computational efficiency, consistent with practices in related studies [21].
@@
where the exponents $H(q)$, $H^{+}(q)$, and $H^{-}(q)$ are the generalized Hurst exponents. They are determined from the slope of a log-log plot of the fluctuation function versus segment size $s$. Our analysis focuses on the case of $q=2$. This choice is standard in the econophysics literature for analyzing volatility persistence, as it directly relates to the second moment (variance) of the fluctuations and provides a measure analogous to the classical Hurst exponent for long-range correlations [15, 25]. Here, $H(2)>0.5$ suggests persistent cross-correlations and $H(2)<0.5$ suggests anti-persistent cross-correlations.
@@ 3.2.1 Core Engine: Lee Oscillator with Retrograde Signaling (LORS)
The decision to replace a static activation function with a dynamic oscillator represents a fundamental paradigm shift grounded in both econophysics and neuroscience. Classical neural networks, with their simple neuron models, have been criticized for being far simpler than their biological counterparts. Modern neuroscience reveals that the brain does not operate on a simple feed-forward firing of static units; rather, it functions on a substrate of continuous and chaotic oscillations known as brainwaves [22]. It is this underlying oscillatory dynamic that gives rise to complex cognitive functions like memory and perception. Inspired by this, we consider that financial markets, as complex adaptive systems driven by collective human behavior, are more faithfully modeled as a field of interacting oscillators than as a system mapped by static functions.
The Lee oscillator, with its design rooted in emulating biologically plausible neural dynamics such as Progressive Memory Recall [26], provides the ideal computational primitive for this paradigm. It allows a system to perform gradual feedback and self-correction when faced with incomplete or noisy inputs through transient chaotic behavior, a mechanism strikingly similar to how market participants adapt to new information. Specifically, this study employs an advanced variant: the Lee Oscillator with Retrograde Signaling. This model enhances the original design [23] by incorporating retrograde signaling mechanisms observed in neuroscience [27, 28], further boosting its biological plausibility. The neural architecture of the LORS is depicted in Figure 2.
@@
A key aspect of our work is the utilization of a diverse set of ten parameterized Lee oscillators. The initial eight types, rigorously derived from systematic studies of the oscillator’s bifurcation behavior [10], already provide a broad range of dynamics. These include simple bifurcations (e.g., T1, T4), dense chaotic regions (T2, T3), and more complex structures with periodic windows (T5, T7). To better capture the specific topological features inherent in financial market dynamics, we extend this set through a systematic exploration of the parameter space. This leads to the design of two additional configurations, T9 and T10, which were empirically identified through parameter space exploration to better capture the unique dynamics of financial markets.
@@ 3.2.3 Stage 2: Final Activation via Maximum Response Selection
Once the library of ten meta-activation functions is generated for a given input $x$ (producing the vector $\mathbf{A}(x)$ as per Algorithm 2), the final step is to produce a single activation output. For this, we adopt the Maximum Response Selection (Max-Select) strategy. This choice is not arbitrary but is deliberately grounded in the Winner-Takes-All (WTA) principle, a computational mechanism with deep roots in theoretical neuroscience and continued relevance in modern cognitive science. The WTA concept, foundational to the theory of efficient neural coding [29], posits that from a group of competing processing units, only the one with the strongest response should determine the final output. This principle is considered a cornerstone of models explaining decision-making and perception [30].
@@ 4.1 Data Description and Feature Generation
We employ the daily realized volatility (RV) as the proxy of the true latent volatility, as it is nearly unbiased and efficient [31]. For a given trading day $t$ with $M$ intraday returns, RV is formally defined as:
@@
The second input series is the volatility increment ($v_{t}$), which captures the dynamics of volatility changes. It is constructed using the realized bipower variation ($BPV_{t}$), a measure known for its robustness to price jumps [32]. For a day $t$ with $M$ intraday returns $r_{t,j}$, BPV is defined as:
@@ 4.2 Experimental Setup
For the training process, all models are optimized by minimizing the Mean Squared Error (MSE) loss function using the Adam optimizer. The MSE is chosen as the training objective due to its stability and its property of heavily penalizing large prediction errors. For hyperparameter tuning on the validation set, we adopt a holistic evaluation approach. While all four performance metrics - MSE, Mean Absolute Error (MAE), the Coefficient of Determination (R²), and the Quasi-Likelihood (QLIKE) - are considered to ensure a well-rounded performance, we place a particular emphasis on the QLIKE loss [33]. As a robust loss function for comparing volatility models under noisy conditions, QLIKE serves as a key criterion for final model selection, ensuring that our chosen configurations are both accurate and superior from a risk-management perspective. Key hyperparameters, such as the learning rate (with a primary value of 0.001) and number of hidden neurons, are tuned via the grid search, with a fixed look-back window of 60 days and a batch size of 64.
To ensure the robustness of our findings, we apply the Model Confidence Set (MCS) procedure [34] to identify an optimal set of models with predictive advantages, considering that multiple models can be considered the best with the equal predictive accuracy. The MCS test is applied to a pool of models comprising the benchmark models and their variants enhanced with either the standard MF-ADCCA or our proposed OSW-MF-ADCCA features. This comparison allows us to statistically validate the effectiveness of our feature engineering approach before integrating it into the final FCOC framework. At a 25% significance level, we conduct the procedure with 10,000 bootstrap resamples and a block length of 30 to generate the MCS p-values.

### v2

@@ 1 Introduction
The forecasting of financial market volatility, as the second moment of asset returns, is a cornerstone of modern finance, crucial for risk management, option pricing, and portfolio allocation [taylor2011asset]. While deep learning methods have shown immense potential in this domain due to their powerful nonlinear pattern recognition capabilities [kim2018forecasting, liu2019novel], their full potential remains constrained by two fundamental and often-overlooked bottlenecks in the standard modeling pipeline.
The first bottleneck is feature fidelity. Financial time series, as typical complex systems, possess an intrinsic structure that traditional statistical features cannot fully capture. The Fractal Market Hypothesis (FMH) from econophysics [peters1994fractal] points out that long-range memory and multifractal properties are key dynamical characteristics of the market. This makes fractal analysis a powerful and theoretically grounded tool for our investigation. However, the application of these advanced tools faces significant limitations. First, their analytical framework is often limited to a single time series, largely failing to consider the cross-asset risk transmission mechanisms, thus ignoring the asymmetric cross-correlation phenomenon that is at the core of systemic risk [longin2001extreme, ang2002asymmetric, ding2011asymmetric]. Second, the standard implementation of these methods has inherent instability, is prone to introducing spurious noise, and thus harms the fidelity of the extracted features [bashan2008comparison]. These two problems together lead to a situation where the input information the model relies on is, at its source, a distorted representation of the market’s true state.
The second, and more fundamental, bottleneck lies in model responsiveness. The core processing units widely adopted in deep learning models—static activation functions such as ReLU—reveal their inherent design limitations when processing highly dynamic and non-stationary financial signals. This limitation stems from a complexity mismatch [wang2021chaotic]: a simple, fixed-logic processor is used to analyze an essentially chaotic and dynamic complex signal. This bottleneck means that even with perfect input features, the model’s rigid internal processing mechanism cannot provide an effective dynamic response, ultimately leading to information decay and loss during internal transmission [lee2019chaotic].
@@ 2.1 Deep Learning Models for Financial Time Series Forecasting
Modern financial forecasting methods have widely adopted deep learning architectures. Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) [kim2018forecasting, liu2019novel] and Gated Recurrent Units (GRU) [chen2019forecasting], are commonly used for their ability to handle temporal dependencies. In recent years, more advanced architectures such as the Transformer [vaswani2017attention] and state-space models like Mamba have also been explored. For instance, Lu and Xu [lu2024trnn] proposed an efficient Time-series Recurrent Neural Network (TRNN) to improve training efficiency. Chen et al. [chen2024improved] developed a hybrid model combining a Temporal Convolutional Network (TCN) with BiGRU for new energy stock index forecasting. While these works have made significant progress in model architecture, they almost universally rely on traditional static activation functions, thus failing to fundamentally address the internal model responsiveness problem.
@@ 2.2 Multifractal Analysis of Financial Time Series
To overcome the informational limitations of raw time series, a more theoretically profound approach originates from econophysics. This discipline aims to uncover the intrinsic physical laws of the market, with the multifractal nature of financial time series being a key insight that reveals the non-uniformity of volatility across different time scales [kantelhardt2002multifractal].
In response, econophysics has developed an evolving toolkit of methods, progressing from Detrended Fluctuation Analysis (DFA) [peters1994fractal] to Multifractal DFA (MF-DFA) [kantelhardt2002multifractal]. A significant advance came with the recognition of market asymmetry, leading to Asymmetric MF-DFA (A-MFDFA) [lee2017asymmetric, lee2018asymmetric]. However, these methods were long confined to single-asset analysis, overlooking the systemic interactions of the market as a whole. In fact, asymmetric cross-correlations between different assets are a key driver of systemic market risk, especially during downturns [longin2001extreme, ang2002asymmetric, ding2011asymmetric]. Recent work has continued to validate this, with Yu et al. [yu2022novel] successfully combining multifractal analysis with GRU networks and Wang and Lee [wang2023stock] applying modified MF-ADCCA to forecast stock index volatility. Although subsequent research began to address cross-correlations with more complex techniques, they commonly inherited a persistent technical flaw from earlier methods: the use of non-overlapping segmentation, which has been shown to introduce spurious fluctuations and compromise the stability of fractal measurements [bashan2008comparison]. Our Fractal Feature Corrector (FFC) is designed to solve both of these historical problems simultaneously. By using Multifractal Asymmetric Detrended Cross-Correlation Analysis (MF-ADCCA) [cao2014detrended] as its core to capture systemic interactions, and employing a robust Overlapping Sliding Window (OSW) implementation to correct for instability [tang2019research], the FFC aims to provide a more reliable and higher-fidelity set of complexity features than previously available.
@@ 2.3 Dynamic and Chaotic Activation Functions
The concept of enhancing neural networks with internal dynamic components originates from computational neuroscience. Unlike standard neural networks, the biological brain is believed to operate based on continuous and chaotic neural oscillations [freeman2000neurodynamics]. Inspired by this, academia has begun to explore replacing static activation functions with dynamic systems [wang2021chaotic]. This direction also aligns with cutting-edge paradigms in econophysics, which posit that financial markets can be viewed as a dynamic energy field where assets behave like interacting, chaotic oscillators. From this perspective, an oscillator with excitatory and inhibitory mechanisms is not merely a computational tool but a physical analogy for the fundamental push-pull dynamics of market forces.
Therefore, our adoption of the Chaotic Oscillation Component (COC) is dually motivated: computationally, it solves the complexity mismatch problem; physically, it provides a more plausible model for the phenomena we aim to forecast. Our COC is built upon the foundational work on the Lee oscillator [lee2004transient]. While generative models like the VAR-VAE proposed by Leushuis [leushuis2025probabilistic] have started to incorporate probabilistic dynamics in latent spaces, our work focuses on a different, complementary goal: embedding dynamics directly into the neuron’s fundamental activation process. Our core contribution in this area is not only the systematic application of this concept to financial volatility forecasting but also the engineering of two novel oscillator configurations (T9 and T10) through parameter space exploration, specifically to better capture the unique dynamics of financial markets, such as abrupt regime switching.
@@ 3.1 Fractal Feature Corrector (FFC)
This process yields $N_{seg}=\lfloor(N-s)/s_{step}\rfloor+1$ overlapping segments. In this work, an overlap ratio of $\rho=1/3$ is used, a choice that strikes a balance between enhancing statistical stability and maintaining computational efficiency, consistent with practices in related studies [tang2019research].
@@
where the exponents $H(q)$, $H^{+}(q)$, and $H^{-}(q)$ are the generalized Hurst exponents. They are determined from the slope of a log-log plot of the fluctuation function versus segment size $s$. Our analysis focuses on the case of $q=2$. This choice is standard in the econophysics literature for analyzing volatility persistence, as it directly relates to the second moment (variance) of the fluctuations and provides a measure analogous to the classical Hurst exponent for long-range correlations [kantelhardt2002multifractal, podobnik2008detrended]. Here, $H(2)>0.5$ suggests persistent cross-correlations and $H(2)<0.5$ suggests anti-persistent cross-correlations.
@@ 3.2.1 Core Engine: Lee Oscillator with Retrograde Signaling (LORS)
The decision to replace a static activation function with a dynamic oscillator represents a fundamental paradigm shift grounded in both econophysics and neuroscience. Classical neural networks, with their simple neuron models, have been criticized for being far simpler than their biological counterparts. Modern neuroscience reveals that the brain does not operate on a simple feed-forward firing of static units; rather, it functions on a substrate of continuous and chaotic oscillations known as brainwaves [freeman2000neurodynamics]. It is this underlying oscillatory dynamic that gives rise to complex cognitive functions like memory and perception. Inspired by this, we consider that financial markets, as complex adaptive systems driven by collective human behavior, are more faithfully modeled as a field of interacting oscillators than as a system mapped by static functions.
The Lee oscillator, with its design rooted in emulating biologically plausible neural dynamics such as Progressive Memory Recall [lee2006progressive], provides the ideal computational primitive for this paradigm. It allows a system to perform gradual feedback and self-correction when faced with incomplete or noisy inputs through transient chaotic behavior, a mechanism strikingly similar to how market participants adapt to new information. Specifically, this study employs an advanced variant: the Lee Oscillator with Retrograde Signaling. This model enhances the original design [lee2004transient] by incorporating retrograde signaling mechanisms observed in neuroscience [levitan2002neuron, wong2008wind], further boosting its biological plausibility. The neural architecture of the LORS is depicted in Figure 2.
@@
A key aspect of our work is the utilization of a diverse set of ten parameterized Lee oscillators. The initial eight types, rigorously derived from systematic studies of the oscillator’s bifurcation behavior [lee2019chaotic], already provide a broad range of dynamics. These include simple bifurcations (e.g., T1, T4), dense chaotic regions (T2, T3), and more complex structures with periodic windows (T5, T7). To better capture the specific topological features inherent in financial market dynamics, we extend this set through a systematic exploration of the parameter space. This leads to the design of two additional configurations, T9 and T10, which were empirically identified through parameter space exploration to better capture the unique dynamics of financial markets.
@@ 3.2.3 Stage 2: Final Activation via Maximum Response Selection
Once the library of ten meta-activation functions is generated for a given input $x$ (producing the vector $\mathbf{A}(x)$ as per Algorithm 2), the final step is to produce a single activation output. For this, we adopt the Maximum Response Selection (Max-Select) strategy. This choice is not arbitrary but is deliberately grounded in the Winner-Takes-All (WTA) principle, a computational mechanism with deep roots in theoretical neuroscience and continued relevance in modern cognitive science. The WTA concept, foundational to the theory of efficient neural coding [barlow1972single], posits that from a group of competing processing units, only the one with the strongest response should determine the final output. This principle is considered a cornerstone of models explaining decision-making and perception [heathcote2022winner].
@@ 4.1 Data Description and Feature Generation
We employ the daily realized volatility (RV) as the proxy of the true latent volatility, as it is nearly unbiased and efficient [andersen2003modeling]. For a given trading day $t$ with $M$ intraday returns, RV is formally defined as:
@@
The second input series is the volatility increment ($v_{t}$), which captures the dynamics of volatility changes. It is constructed using the realized bipower variation ($BPV_{t}$), a measure known for its robustness to price jumps [barndorff2004power]. For a day $t$ with $M$ intraday returns $r_{t,j}$, BPV is defined as:
@@ 4.2 Experimental Setup
For the training process, all models are optimized by minimizing the Mean Squared Error (MSE) loss function using the Adam optimizer. The MSE is chosen as the training objective due to its stability and its property of heavily penalizing large prediction errors. For hyperparameter tuning on the validation set, we adopt a holistic evaluation approach. While all four performance metrics - MSE, Mean Absolute Error (MAE), the Coefficient of Determination (R²), and the Quasi-Likelihood (QLIKE) - are considered to ensure a well-rounded performance, we place a particular emphasis on the QLIKE loss [patton2011volatility]. As a robust loss function for comparing volatility models under noisy conditions, QLIKE serves as a key criterion for final model selection, ensuring that our chosen configurations are both accurate and superior from a risk-management perspective. Key hyperparameters, such as the learning rate (with a primary value of 0.001) and number of hidden neurons, are tuned via the grid search, with a fixed look-back window of 60 days and a batch size of 64.
To ensure the robustness of our findings, we apply the Model Confidence Set (MCS) procedure [hansen2011model] to identify an optimal set of models with predictive advantages, considering that multiple models can be considered the best with the equal predictive accuracy. The MCS test is applied to a pool of models comprising the benchmark models and their variants enhanced with either the standard MF-ADCCA or our proposed OSW-MF-ADCCA features. This comparison allows us to statistically validate the effectiveness of our feature engineering approach before integrating it into the final FCOC framework. At a 25% significance level, we conduct the procedure with 10,000 bootstrap resamples and a block length of 30 to generate the MCS p-values.